OpenStudy (anonymous):

Solve the equation.... 4(2x-3)=6-(3-2x)

OpenStudy (anonymous):

8x - 12 = 6 -3 + 2x 6x = 15 x = 15 / 6 = 5/2

OpenStudy (anonymous):

distribute the equation (remove the parentheses) 8x-12=6-3+2x 8x-12=3+2x 6x=15 x=15/6=5/2